Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

DURC 50m Open

Distribution

Dr. Trevor West DUCAC Chairman
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer

Summary

Venue Fassaroe Sporting Club
Date July 22
Discipline 50m Prone Rifle

DURC Attendees

Steven Watterson  Ray Reilly
Peter Byrne  Alan Smyth
Mark Dennehy  John Keeney

Adrian Costigan

Scores

Steven Watterson 585
Ray Reilly 582
Peter Byrne 577
Alan Smyth 572
Mark Dennehy 568
John Keeney 566
Adrian Costigan 545

Prizes Won

Steven Watterson 1st place overall, 2nd place class A, 2nd place Rapid-Fire
Ray Reilly 2nd place overall,3rd place class A, 1st place rapid-fire
Adrian Costigan 2nd place class C
Alan Smyth 1st place class D
Peter Byrne 2nd place class D

Report

Steven Watterson won the Dublin University Rifle Club Open today, beating Olympian Alan Lewis with a score of 585. Steven, an Isle of Man native, has been shooting for many years in the Isle of Man and Scotland. Currently the captain of the Dublin University Rifle Club, he has just returned from the Island Games in the Isle of Man where he won two silver medals in Air Rifle and 50m three-position Rifle, as well as a bronze medal.

Steven won a custom rifle barrel kindly donated by Shilen ltd. Second place went to Ray Reilly and third to Tony Cogan who is travelling to Zagreb later this week to represent Ireland in the European Championships.

The Rapid-Fire competition was won by Raymond Reilly, with an impressive 97/100. Second place was decided by a shoot-off between Steven Watterson and Alan Lewis, and was won by Steven Watterson.

Misé le meas,

Mark Dennehy, DURC Secretary